Circular No. 7029

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1987S, 1988af, 1988ag, 1996co Jean Mueller reports her discoveries of four apparent supernovae on

photographs taken with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey:

SN  UT  α_{2000}  δ_{2000}  Mag.  Offset
1987S 1987 Aug. 28 21^{h}49^{m}45.^{s}05 +12^{o}12′54′′.4 18  3′′ E, 10′′ N
1988af 1988 Nov. 19 8 17 24.36 +45 38 24.9 17.5 5′′.8 E, 7′′.6 S
1988ag 1988 Dec. 11 9 56 45.85 + 6 56 22.5 18  8′′.7 E, 7′′ N
1996co 1996 Dec. 20 10 18 57.48 +62 15 12.1 18 5′′.2 W, 9′′.2 S

SN 1987S also appears at mag ∼ 19 on a sky-survey plate taken by J. Phinney and Mueller on 1987 Sept. 17; no object appears at the position of SN 1987S on the Digitized Sky Survey or on a IIIa-F sky-survey plate taken on 1990 Oct. 22 by C. Brewer and J. D. Mendenhall. SN 1988af in MCG +08-15-057, discovered on a plate taken by P. Moniot, also appears at mag ∼ 20 on another survey plate taken on 1989 Feb. 27 by Moniot and Brewer; there is no object at the position of SN 1988af on the Digitized Sky Survey or on a sky-survey plate taken on 1990 Mar. 18 by Mueller. SN 1988ag, discovered on a plate taken by Brewer and Mueller, also appears at mag ∼ 18.5 on a 0.46-m Palomar Schmidt telescope film taken by E. M. and C. S. Shoemaker (identified by B. Skiff and C. S. Shoemaker) on 1989 Jan. 9; there is no object at the position of SN 1988ag on the Digitized Sky Survey or on a IIIa-F sky-survey plate taken on 1997 Mar. 9 by Mueller and Mendenhall. SN 1996co, discovered on a plate taken by K. Rykoski and Mendenhall, also appears at mag ∼ 20 on a sky-survey plate taken on 1997 Jan. 11 by Mueller. There is nothing at the position of SN 1996co on a sky-survey plate taken on 1997 May 6 or on CCD frames taken by R. Gal with the Palomar 1.5-m telescope on 1997 May 1.

XTE J1946+274 R. M. Hjellming, National Radio Astronomy Observatory (NRAO);

and A. J. Mioduszewski, Joint Institute for VLBI in Europe and NRAO, report: “Continued Very Large Array observations of the error region of XTE 1946+274 (IAUC 7014, 7016) on Oct. 12 and 13 have failed to confirm the possible radio counterpart we reported on IAUC 7027. Reanalysis of the Oct. 7.02 UT radio data at 1.49 GHz indicates it is possible that the reported radio source was a spurious artifact due to imperfect CLEAN sidelobe subtraction for another strong radio source in the field.”

1998 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green